40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9945     Entry time: Tue May 13 03:30:10 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Locking activities - no progress 

[Jenne, Rana, EricQ]

We tried a few times to engage the AO path while holding CARM on sqrtTrans and DARM on ALS, but failed every time.  Since we cannot stably hold lock at arm powers of 1, even though we were able to do so early last week, we think that we have a problem (obviously).  One noticeable thing is that while held with ALS, the Xarm transmission fluctuates almost full power.

As we were seeing late last week, the Xarm IR transmission while held with ALS was fluctuating wildly, whether we were locked on individual arms or on CARM/DARM. 

Tonight we took some out of loop spectra, with different HEPA settings, to see how that affected things.  It looks like HEPA at the nominal ~20% is okay, but anything higher than that starts to affect the Xarm beatnote sensing, while it mostly leaves the Yarm beatnote sensing okay.  Perhaps this is something that isn't tightened enough in the Xarm path, or something on a skinny floppy mount that needs to be more secure.

I am still a little confused though, why we don't see large power fluctuations in *both* arms while using ALS to control CARM/DARM.  Why are we not seeing this Xarm noise being fed back into the Yarm, through either the ETM via DARM, or common stuff via CARM actuating on MC2.

Note that the change at high frequency is because I switched from using non-DQ channels to DQ channels, so that's not anything to pay attention to.  The noise reduction we see is below about 20Hz.

ALS_outofloop.pdf


Rana pointed out that our POPDC level was very small.  We don't have screens for them, but the DC PDs have the same analog whitening as the RF PD signals do.  I changed C1:LSC-POPDC_WhiteGain.VAL from 0 to 11.  Now our POPDC while locked on sidebands is about 8,000 counts.

We also swapped the cables between the SR785 and the CM board around.  Now channel 2 of the 785 goes to TP2A, channel 1 goes to TP2B, and the source goes to EXCB.  This is so that we can break the AO loop with the disable switch just after the slow/fast split, and look at the transfer function before we close the loop. 

ELOG V3.1.3-